The Sterling Tire Companys income statement for 20X1 is as follows STERLING TIRE COMPANY For the Year Ended Decenber 31, 28x1 Sales (21,800 tires at $62 each) Variable costs (21,038 tires at S31) Fixed costs Earnings before interest and taxes (EBIT) Interest expense Earnings before taxes (EBT) Income tax expense (20%) Earnings after taxes (EAT) 5 1,382,098 651,808 241,890 5,500 eBook 5 190, 59 38,108 5 152,400 e. Compute the degree of operating leverage. (Round your answer to 2 decimel places. Degree of operating leverage b. Compute the degree of financial leverage. (Round your answer to 2 decimal places.) Degree of financial leverage c. Compute the degree of combined leverage. (Round your answer to 2 decimal places.) Degree of combined leverage d. Compute the breek-even point in units. (Round your answer to the nearest whole number.) Break-even point units

The Sterling Tire Company's income statement for 20X1 is as follows 

STERLING TIRE COMPANY For the Year Ended Decenber 31, 28x1 


Sales (21,800 tires at $62 each)  $1,302,000

Variable costs (21,038 tires at S31) 651,000

Fixed costs                               410,000

Earnings before interest and taxes (EBIT)  241,000

Interest expense                       50,500

Earnings before taxes (EBT)      $190,500

Income tax expense (20%)         38,100

Earnings after taxes (EAT)       $152,400 


a. Compute the degree of operating leverage. 

b. Compute the degree of financial leverage.

c. Compute the degree of combined leverage

d. Compute the breek-even point in units.
Answer #2

SOLUTION :


a.


Degree of operating leverage (DOL)

= (Sales - Variable costs) / EBIT

= (1302000 - 651000) / 241000

= 651000 / 241000

= 2.70124

= 2.70 (ANSWER)


b.


Degree of financial leverage (DFL)

= EBIT / EBT 

= 241000 / 190500

= 1.2651

= 1.27 (ANSWER).


c.


Degree of combined leverage 

= Contribution margin / EBT  

= (1302000 - 651000)  / 190500

= 3.4173

= 3.42  (ANSWER).


d.


Break-even Point  in units = Fixed costs / Unit contribution margin

= 410000 / (62 - 31)

= 13225.81

= 13226 tires (ANSWER).


answered by: Tulsiram Garg
Answer #1

Quantity = 21,000

Price = $62

Variable Cost = $30

Fixed Cost = $400,000

Interest expense= $50,000

a

Degree of operating leverage = (sales − variable costs) / (sales − variable costs − fixed costs)

Degree of operating leverage=((62-30)*21000)/((62-30)*21000-400000)

Degree of operating leverage=2.47

b

Degree of Financial Leverage =EBIT/(EBIT-Interest Expense)

=241000/(241000-50000)

Degree of Financial Leverage=1.26

C

Degree of Combined Leverage

Degree of Combined Leverage = (sales − variable costs) / (sales − variable costs − fixed costs-Interest Expense)

=((62-30)*21000)/((62-30)*21000-400000-50000)

Degree of Combined Leverage=3.03

          d

Break even Point

Degree of operating leverage =fixed costs/ (sales − variable costs)

=400000/(62-30)

Break even Point=12500 Units